What mass of KBr is needed to make 63.9 ml of a 0.351 M KBr solution?

_______ G
1
Expert's answer
2020-05-20T13:47:10-0400

m(KBr)=C(KBr)*V*Mr(KBr)=0.351*0.0639*119=2,669 g